# RedundancyLens: Revealing and Exploiting Visual Token Processing Redundancy for Efficient Decoder-Only MLLMs

### v2

@@ Abstract
Hongliang Li1, Jiaxin Zhang1, Wenhui Liao1, Dezhi Peng1, Kai Ding2, Lianwen Jin1thanks: Corresponding author. 1South China University of Technology   2Intsig Information Co., Ltd. eehongliangli@mail.scut.edu.cn, eejxzhang@gmail.com, eelwh@mail.scut.edu.cn pengdzscut@foxmail.com, danny_ding@intsig.net, eelwjin@scut.edu.cn
@@ 2.1 MLLM Architectures
The cross-attention-based architecture offers greater efficiency in handling long visual token sequences, gaining increasing attention as an alternative to decoder-only architecture. These architectures introduce additional cross-attention layers within the LLM to integrate visual information by applying cross-attention to visual tokens, thereby eliminating the need for the entire LLM to process them. Flamingo Alayrac et al. (2022) is a prominent early work in this area, using a perceiver resampler to downsample the vision encoder’s features before feeding them into the LLM via gated cross-attention layers. Llama 3-V Llama-Team (2024) adopts a similar structure but removes the perceiver module. EVLM Chen et al. (2024a) utilizes hierarchical ViT features and a mixture of experts to enhance performance. mPLUG-Owl3 Ye et al. (2024a) incorporates cross-attention mechanisms in parallel with self-attention layers instead of adding additional cross-attention layers. EE-MLLM Ma et al. (2024) modifies the original self-attention mechanism into a composite attention mechanism. Meanwhile, NVLM Dai et al. (2024) introduces a hybrid architecture that uses the LLM’s self-attention layers to process thumbnail image tokens while employing cross-attention to capture finer image details.
@@ 3.1.1 Probe-Activated Dynamic FFN
Inspired by MoE Shazeer et al. (2016), we reduce FFN computations for visual tokens by structurally activating only a subset of FFN parameters. However, we cannot directly adopt MoE, as it requires training a router that dynamically selects which parameters to activate. To achieve this without additional training, we propose the Probe-Activated Dynamic FFN.
@@
Figure 2 (a) provides a more intuitive illustration of the computation process in the Probe-Activated Dynamic FFN, with activation functions and biases omitted for simplicity. It is important to note that this process applies only to visual tokens, while the FFN for text tokens remains unchanged. Some MLLMs modify the vanilla FFN, such as by adding gating mechanisms Wang et al. (2024a); Chen et al. (2024c), yet our method can still be directly applied in these cases.
@@ 3.2 Layer Ranking Algorithm
Given the number of layers requiring reduction, denoted by $L_{r}$ (where $0\leq L_{r}\leq L$, and $L$ is the total number of layers), the goal is to select the $L_{r}$-layer combination with the highest redundancy. To this end, we construct a compact validation set and use the performance variations of the MLLM on it to estimate redundancy. Since exhaustively evaluating all possible layer combinations for each value of $L_{r}$ is computationally infeasible, we propose a search algorithm that ranks each layer, as detailed in Algorithm 1. For a given $L_{r}$, the layers with the highest ranks are prioritized for reduction.
@@
For each MLLM, the algorithm is applied separately to FFN and attention reductions, running the process twice. In our experiments, we observe that the last few layers of MLLMs tend to exhibit greater redundancy, making them a priority for reduction. To reduce the number of evaluations, we limit the ranking algorithm’s search to the first $L-L_{p}$ layers. The last $L_{p}$ layers are ranked in descending order of their position, starting from the last layer.
@@ 4.1 Datasets
To construct the validation set for Layer Ranking Algorithm, we randomly sample 750 instances from the full evaluation dataset collected in Liu et al. (2024c), 200 instances from the validation set of DocVQA Mathew et al. (2021), 200 from the validation set of InfoVQA Mathew et al. (2022), 200 from the validation set of ChartQA Masry et al. (2022), and 1,000 from MMBench-DEV-EN-V11 Liu et al. (2024b). To avoid overlap with the test set, for the 750 instances sampled from the evaluation dataset in Liu et al. (2024c), we remove any data that also appears in OCRBench Liu et al. (2024c) and exclude samples from TextVQA Singh et al. (2019), DocVQA Mathew et al. (2021), InfoVQA Mathew et al. (2022), and ChartQA Masry et al. (2022). For evaluation, we conduct experiments on eight widely used benchmarks: OCRBench Liu et al. (2024c), DocVQA Mathew et al. (2021), InfoVQA Mathew et al. (2022), ChartQA Masry et al. (2022), TextVQA Singh et al. (2019), MME Fu et al. (2024), MMStar Chen et al. (2024b), and HallusionBench Guan et al. (2024). For TextVQA Singh et al. (2019), following previous work Chen et al. (2024c); Wang et al. (2024a), we use its validation set. We assess model performance using the standard metrics provided by each benchmark.
@@ 4.2 Implementation Details
Our experiments are conducted on NVIDIA A100 GPUs using VLMEvalKit Duan et al. (2024), a framework for evaluating large multi-modality models on diverse multimodal benchmarks. We evaluate state-of-the-art MLLMs, including InternVL2-8B Chen et al. (2024c), Qwen2-VL-7B Wang et al. (2024a), MiniCPM-V 2.6 Yao et al. (2024), and LLaVA-OneVision Li et al. (2024a). In Hollow Attention, the attention range $R_{A}$ for visual tokens is set to 256, typically representing the number of tokens in one sub-image. In Probe-Activated Dynamic FFN, the number of randomly sampled visual tokens $M$ is set to 10% of the total visual tokens per sample, while the number of activated parameters $K$ is set to 20% of the original parameter count. During the search process, $M$ is set to 100% of the total visual tokens per sample to minimize fluctuations caused by uncertainty.
@@ 4.4 Comparison with Training-free MLLM Inference Acceleration Methods
Figure 6: Ablation studies on (a) the attention range in Hollow Attention and (b) the proportion of activated parameters in Probe-Activated Dynamic FFN, evaluated on ChartQA by InternVL2-8B. The x-axis represents the percentage of layers with reductions applied, and the y-axis indicates model performance on the benchmark metric.
@@
4.5 Ablation Studies
Ablation Studies on the Extent of Reduction in Self-Attention and FFN. To comprehensively evaluate the impact of reducing self-attention and FFN operations, we conduct ablation studies focusing on two factors: the proportion of activated parameters in the Probe-Activated Dynamic FFN and the attention range in the Hollow Attention, as illustrated in Figure 6. The results show that as the extent of reduction decreases (i.e., as the proportion of activated parameters or the attention range increases), more layers can be reduced without significantly impacting model performance. These two hyperparameters are selected by trading off efficiency and effectiveness.
Ablation Study of Layer Ranking Strategies. We compared three layer ranking strategies: (1) Position-based Strategy, which assigns the highest rank to the last layer and progressively decreases the rank toward the first layers; (2) Search-only strategy, which relies solely on Algorithm 1, with no layers pre-assigned ranks; and (3) Hybrid strategy, where the last $L_{p}$ layers are pre-assigned the highest rank, with the remaining layers ranked by Algorithm 1. As shown in Figure 7, when reduction is applied to only a few layers, the position-based strategy outperforms the search-only strategy, indicating that later layers tend to exhibit higher redundancy in visual token processing. Additionally, the limited size of our validation set may not fully capture the true behavior of the models. As the number of reduced layers increases, the search-only strategy begins to yield better results. Therefore, we adopt the hybrid strategy, which combines the position-based strategy with the search-only strategy, to achieve better performance and reduce the number of evaluations required by Algorithm 1.
@@ 5 Conclusion
Appendix A Evaluation Metrics for Each Benchmark
OCRBench Liu et al. (2024c) uses the number of correctly generated answers as its evaluation metric. DocVQA Mathew et al. (2021) and InfoVQA Mathew et al. (2022) use Average Normalized Levenshtein Similarity (ANLS) and are evaluated on their respective official websites. ChartQA Masry et al. (2022) measures performance with relaxed accuracy, while TextVQA Singh et al. (2019) relies on VQA accuracy Goyal et al. (2017). MME Fu et al. (2024) reports the sum of perception and cognition scores. MMStar Chen et al. (2024b) evaluates models based on overall accuracy. HallusionBench Guan et al. (2024) reports the average of Question Pair Accuracy, Figure Accuracy, and Overall Accuracy. All datasets, except DocVQA and InfoVQA, are evaluated using the VLMEvalKit Duan et al. (2024).
The validation set for Layer Ranking Algorithm consists of multiple subsets from different datasets. The total evaluation score is obtained by summing the scores across all subsets. For each subset, we calculate the difference between the evaluation metric of the reduced model and the original model. If the difference is negative, meaning the reduced model performs worse, it is multiplied by a penalty coefficient $\alpha>1$ to amplify the impact of the performance drop. This approach ensures that performance drops are further penalized, guiding the reduction strategy toward solutions with more stable overall performance across the entire validation set. The penalty coefficient $\alpha$ is set to 2.

### v3

@@ Abstract
Hongliang Li1, Jiaxin Zhang1, Wenhui Liao1, Dezhi Peng1, Kai Ding2,3, Lianwen Jin11footnotemark: 11,3 1South China University of Technology  2Intsig Information Co., Ltd. 3INTSIG-SCUT Joint Lab on Document Analysis and Recognition eehongliangli@mail.scut.edu.cn, eejxzhang@gmail.com, eelwh@mail.scut.edu.cn pengdzscut@foxmail.com, danny_ding@intsig.net, eelwjin@scut.edu.cn
@@ 2.1 MLLM Architectures
The cross-attention-based architecture offers greater efficiency in handling long visual token sequences, gaining increasing attention as an alternative to the decoder-only architecture. These architectures introduce additional cross-attention layers within the LLM to integrate visual information by applying cross-attention to visual tokens, thereby eliminating the need for the entire LLM to process them. Flamingo Alayrac et al. (2022) is a prominent early work in this area, using a perceiver resampler to downsample the vision encoder’s features before feeding them into the LLM via gated cross-attention layers. Llama 3-V Llama-Team (2024) adopts a similar structure but removes the perceiver module. EVLM Chen et al. (2024a) utilizes hierarchical ViT features and a mixture of experts to enhance performance. mPLUG-Owl3 Ye et al. (2024a) incorporates cross-attention mechanisms in parallel with self-attention layers instead of adding additional cross-attention layers. EE-MLLM Ma et al. (2024) modifies the original self-attention mechanism into a composite attention mechanism. Meanwhile, NVLM Dai et al. (2024) introduces a hybrid architecture that uses the LLM’s self-attention layers to process thumbnail image tokens while employing cross-attention to capture finer image details.
@@ 3.1.1 Probe-Activated Dynamic FFN
Inspired by MoE Shazeer et al. (2016), we reduce FFN computations for visual tokens by structurally activating only a subset of FFN parameters. However, we cannot directly adopt MoE, as it requires training a router that dynamically selects which parameters to activate. To achieve this without additional training, we propose Probe-Activated Dynamic FFN.
@@
Figure 2 (a) provides a more intuitive illustration of the computation process in Probe-Activated Dynamic FFN, with activation functions and biases omitted for simplicity. It is important to note that this process applies only to visual tokens, while the FFN for text tokens remains unchanged. Some MLLMs modify the vanilla FFN, such as by adding gating mechanisms Wang et al. (2024a); Chen et al. (2024c), yet our method can still be directly applied in these cases.
@@ 3.2 Layer Ranking Algorithm
Given the number of layers requiring reduction, denoted by $L_{r}$ (where $0\leq L_{r}\leq L$, and $L$ is the total number of layers), the goal is to select the $L_{r}$-layer combination with the highest redundancy. To this end, we construct a compact validation set and use the performance variations of the MLLM on it to estimate redundancy. Since exhaustively evaluating all possible layer combinations for each value of $L_{r}$ is computationally infeasible, we propose a search algorithm that ranks each layer, as detailed in Algorithm 1. For a given $L_{r}$, the top-ranked $L_{r}$ layers are selected for reduction.
@@
The validation set used in the Layer Ranking Algorithm comprises multiple subsets drawn from different datasets. The overall evaluation metric is computed by summing the scores across all subsets. For each subset, we compute the difference by subtracting the evaluation metric of the original model from that of the reduced model. If this difference is negative—indicating that the reduced model performs worse—it is multiplied by a penalty coefficient $\alpha>1$. This penalization mechanism encourages the search process to prioritize reductions that maintain performance stability.
For each MLLM, Algorithm 1 is applied separately for FFN and attention reductions, with the process run twice. In our experiments, we observe that the last few layers of MLLMs tend to exhibit greater redundancy, making them a priority for reduction. To reduce the number of evaluations, we limit the ranking algorithm’s search to the first $L-L_{p}$ layers. The last $L_{p}$ layers are ranked in descending order of their position, starting from the last layer.
@@ 4.1 Datasets
To construct the validation set for the Layer Ranking Algorithm, we randomly sample 750 instances from the full evaluation dataset collected in Liu et al. (2024c), 1,000 instances from MMBench-DEV-EN-V11 Liu et al. (2024b), and 200 instances each from the validation sets of DocVQA Mathew et al. (2021), InfoVQA Mathew et al. (2022), and ChartQA Masry et al. (2022). To avoid overlap with the test set, when sampling the 750 instances from the full evaluation dataset in Liu et al. (2024c), we exclude any samples that appear in OCRBench Liu et al. (2024c), as well as those from TextVQA Singh et al. (2019), DocVQA Mathew et al. (2021), InfoVQA Mathew et al. (2022), and ChartQA Masry et al. (2022).
For evaluation, we conduct experiments on eight widely used benchmarks: OCRBench Liu et al. (2024c), DocVQA Mathew et al. (2021), InfoVQA Mathew et al. (2022), ChartQA Masry et al. (2022), TextVQA Singh et al. (2019), MME Fu et al. (2024), MMStar Chen et al. (2024b), and HallusionBench Guan et al. (2024). In particular, following prior work Chen et al. (2024c); Wang et al. (2024a), we use the validation set of TextVQA Singh et al. (2019) for evaluation.
4.2 Evaluation Metrics for Each Benchmark
We assess model performance using the standard metrics provided by each benchmark. OCRBench Liu et al. (2024c) uses the number of correctly generated answers as its evaluation metric. DocVQA Mathew et al. (2021) and InfoVQA Mathew et al. (2022) use Average Normalized Levenshtein Similarity (ANLS) and are evaluated on their respective official websites. ChartQA Masry et al. (2022) measures performance with relaxed accuracy, while TextVQA Singh et al. (2019) relies on VQA accuracy Goyal et al. (2017). MME Fu et al. (2024) reports the sum of perception and cognition scores. MMStar Chen et al. (2024b) evaluates models based on overall accuracy. HallusionBench Guan et al. (2024) reports the average of Question Pair Accuracy, Figure Accuracy, and Overall Accuracy. Evaluation for all benchmarks, except DocVQA and InfoVQA, is conducted using VLMEvalKit Duan et al. (2024).
@@ 4.3 Implementation Details
All experiments are conducted on NVIDIA A100 GPUs using VLMEvalKit Duan et al. (2024), a framework for evaluating MLLMs on diverse multimodal benchmarks. We evaluate state-of-the-art MLLMs, including InternVL2-8B Chen et al. (2024c), Qwen2-VL-7B Wang et al. (2024a), MiniCPM-V 2.6 Yao et al. (2024), and LLaVA-OneVision Li et al. (2024a). In Hollow Attention, the attention range $R_{A}$ for visual tokens is set to 256, which typically corresponds to the number of tokens in a single sub-image. In Probe-Activated Dynamic FFN, the number of randomly sampled visual tokens $M$ is set to 10% of the total visual tokens per sample, while the number of activated parameters $K$ is set to 20% of the original parameter count. For the Layer Ranking Algorithm, the penalty coefficient $\alpha$ is set to 2. During the ranking process, $M$ is set to 100% of the total visual tokens per sample to minimize fluctuations caused by uncertainty.
@@ 4.5 Comparison with Training-free MLLM Inference Acceleration Methods
4.6 Ablation Studies
Ablation Studies on the Extent of Reduction in Self-Attention and FFN. To comprehensively evaluate the impact of reducing self-attention and FFN operations, we conduct ablation studies focusing on two factors: the proportion of activated parameters in Probe-Activated Dynamic FFN and the attention range in Hollow Attention, as illustrated in Figure 6. The results show that as the extent of reduction decreases (i.e., as the proportion of activated parameters or the attention range increases), more layers can be reduced without significantly impacting model performance. These two hyperparameters are selected by trading off efficiency and effectiveness.
Figure 6: Ablation studies on (a) the attention range in Hollow Attention and (b) the proportion of activated parameters in Probe-Activated Dynamic FFN, evaluated on ChartQA by InternVL2-8B. The x-axis indicates the percentage of layers with reductions applied, while the y-axis reflects the model’s performance.
Ablation Study of Layer Ranking Strategies.
We compared three layer ranking strategies: (1) Position-based Strategy, which assigns the highest rank to the last layer and progressively decreases the rank toward the first layers; (2) Search-only strategy, which relies solely on Algorithm 1, with no layers pre-assigned ranks; and (3) Hybrid strategy, where the last $L_{p}$ layers are pre-assigned the highest rank, with the remaining layers ranked by Algorithm 1. As shown in Figure 7, when reduction is applied to only a few layers, the position-based strategy outperforms the search-only strategy, indicating that later layers tend to exhibit higher redundancy in visual token processing. Additionally, the limited size of our validation set may not fully capture the true behavior of the models. As the number of reduced layers increases, the search-only strategy begins to yield better results. Therefore, we adopt the hybrid strategy, which combines the position-based strategy with the search-only strategy, to achieve better performance and reduce the number of evaluations required by Algorithm 1.
@@ 5 Conclusion
Acknowledgements
This research is supported in part by the National Natural Science Foundation of China (Grant No.: 62476093, 62441604).